2q+3/3 + 3q-4/6 = q-2/2
Is it -5, 5, 2, -2

Answer by checkley77(12844)

2q+3/3 + 3q-4/6 = q-2/2
2q+1+3q-2/3=q-1
5q-q=-1-1+2/3
4q=-2+2/3
4q=-4/3
q=(-4/3 )/4
q=-4/3*1/4
q=-4/12
q=-1/3 ans.
Is it -5, 5, 2, -2 NO.